# Instability of the Haldane Phase: Roles of Charge Fluctuations and Hund’s Coupling

### v1

@@ Abstract
Introduction.— Understanding the robustness of topological phases in the presence of competing interactions and additional degrees of freedom remains a central challenge in condensed matter physics. Symmetry-protected topological (SPT) phases, exemplified by the Haldane phase [1, 2] in spin-1 chains, exhibit hallmark features including a finite spin gap [3, 4], nonlocal string order [5, 6, 7], and fractionalized edge states [8, 9] protected by discrete symmetries [10, 11, 12, 13]. While these properties have been thoroughly established for idealized spin-only models, their stability in realistic materials featuring significant charge fluctuations and orbital degrees of freedom remains poorly understood and is a highly nontrivial issue.
@@
Model definitions and effective mapping.— We consider two models of coupled one-dimensional (1D) chains with interchain ferromagnetic (FM) Hund’s coupling: a spin-$\frac{1}{2}$ Heisenberg ladder and a half-filled Hubbard ladder. The Heisenberg ladder serves as a spin-only reference system, while the latter incorporates charge degrees of freedom and their fluctuations.
@@
[…]
@@
where $\tilde{\mathbf{S}}_{i}=\mathbf{S}_{1,i}+\mathbf{S}_{2,i}$ is a spin-1 operator on rung $i$, and $L$ is the number of rungs.
@@
whose asymptotic value $|\mathcal{O}_{\mathrm{str}}(-\infty,\infty)|$ serves as the string order parameter. For the spin-1 Heisenberg chain, this parameter takes the value $|\mathcal{O}_{\mathrm{str}}^{S=1}(-\infty,\infty)|\approx 0.3743$ [4, 34].
@@
Figure 2: String order in the Heisenberg ladder. (a) String correlation function $\mathcal{O}_{\mathrm{str}}(|i-j|)$ plotted as a function of distance $|i-j|$ for several values of $J_{\rm H}/J$. The dotted line indicates the string order parameter of the $S=1$ Heisenberg chain. (b) String order parameter $|\mathcal{O}_{\rm str}(-\infty,\infty)|$ as a function of $J_{\rm H}/J$. Inset: $|\mathcal{O}_{\rm str}(-\infty,\infty)|$ plotted against $\sqrt{J_{\rm H}/J}$ in the small $J_{\rm H}/J$ regime; the dashed line shows a fit near $J_{\rm H}/J=0$.
Haldane Phase in the Heisenberg Ladder.— We first investigate how the Heisenberg ladder converges to the effective spin-1 Heisenberg chain [Eq. (3)] as $J_{\rm H}/J$ increases. Figure 2(a) shows the string correlation function $|\mathcal{O}_{\rm str}(i,j)|$ versus distance $|i-j|$ for several values of $J_{\rm H}/J$, computed on a $193\times 2$ ladder. For all $J_{\rm H}/J$ values considered, the correlation function clearly saturates at large distances, confirming robust long-range string order.
In Fig. 2(b), we plot the extracted string order parameter as a function of $J_{\rm H}/J$. As $J_{\rm H}/J$ grows, the string order parameter continuously approaches the known spin-1 Heisenberg chain value, demonstrating that the ladder progressively realizes the effective spin-1 limit. Remarkably, as shown in the inset, the string order parameter follows the scaling form $|\mathcal{O}_{\mathrm{str}}(i,j)|\propto\sqrt{J_{\rm H}/J}$ in the small-$J_{\rm H}/J$ regime, indicating the immediate onset of the Haldane phase for any finite $J_{\rm H}>0$. Consistent with this, the spin gap opens simultaneously with the emergence of string order, as discussed in detail below.
@@
Figure 4: Topological string order in the Hubbard ladder. String correlation function $O_{\text{str}}(i,j)$ versus $|i-j|$ for (a) various $J_{H}/t$ at fixed $U/t=4$ and (b) various $U/t$ at fixed $J_{H}/t=4$. Log-log plots of $O_{\text{str}}(i,j)$ versus $|i-j|$ for (c) various interaction strengths with $U/t=J_{H}/t$, and (d) various $J_{H}/t$ at fixed $U/t=0$. String order parameter as a function of (e) $U/t=J_{H}/t$ and (f) $J_{H}/t$ at fixed $U/t=0$. Dotted lines mark the spin-1 Heisenberg chain value.
Topological Order in the Hubbard Ladder.— We now investigate how charge fluctuations influence topological order in the Hubbard ladder by examining the string correlation function $\mathcal{O}_{\rm str}(i,j)$. Figure 4(a) presents $|\mathcal{O}_{\rm str}(i,j)|$ as a function of distance $|i-j|$ for several values of $J_{\rm H}/t$ at fixed $U/t=4$, while Fig. 4(b) shows similar results for various $U/t$ at fixed $J_{\rm H}/t=4$. In both cases, the string correlation saturates at finite values at large distances, confirming robust long-range order. With increasing $U/t$ or $J_{\rm H}/t$, charge fluctuations are progressively suppressed, driving the system closer to the effective spin-1 regime. Correspondingly, the saturated value, i.e., the string order parameter, approaches the spin-1 Heisenberg chain value. Thus, the Hubbard ladder can realize a stable Haldane phase despite moderate charge fluctuations.
@@
Remarkably, long-range string order appears even at $U/t=0$, driven solely by Hund’s coupling. Figure 4(d) shows log-log plots of the string correlation for various $J_{\rm H}/t$ values at zero on-site repulsion. A clear transition from power-law decay ($J_{\rm H}/t=3$) to finite long-range order ($J_{\rm H}/t\geq 6$) is observed, reflecting strong suppression of charge fluctuations (double occupancy) by Hund’s coupling alone. This insulating phase stabilized purely by Hund’s coupling can thus be viewed as a novel “Hund insulator”. The string order parameter, shown in Fig. 4(f), grows around $J_{\rm H}/t\approx 4.0$, approaching the spin-1 Heisenberg value at large $J_{\rm H}/t$, again following an approximately linear onset.
The resulting phase diagram, summarized in Fig. 3, reveals a clear boundary between the trivial Mott/Hund insulating phase and the SPT Haldane phase, identified explicitly by the string order parameter. Interestingly, the Haldane phase remains stable under substantial charge fluctuations, persisting up to a local charge variance of approximately $\delta n^{2}=2\langle n_{\gamma,i,\uparrow}n_{\gamma,i,\downarrow}\rangle\lesssim 0.4$. Furthermore, the phase boundary approaches $J_{\rm H}=0$ in the $U=\infty$ limit, indicating a smooth crossover from the Hubbard to the Heisenberg ladder in the strongly correlated regime. Although our model setup is different, the qualitative agreement with the topological transition line reported in Ref. [25] supports the broader conclusion that the Haldane phase is generally stable against moderate charge fluctuations.
Figure 5: (a) Spin gap $\Delta_{\rm s}/J$ for the Heisenberg ladder as a function of $J_{\rm H}/J$. The dashed line marks the spin gap of the spin-1 Heisenberg chain [Eq. (3)]. Inset: Enlarged view of the small-$J_{\mathrm{H}}/J$ region. (b) $\Delta_{\rm s}/t$ for the Hubbard ladder as a function of $J_{\rm H}/t$ for several values of $U/t$. (c) $\Delta_{\rm s}/t$ as a function of $J_{\rm H}/t$ at $U=0$. Dashed lines in (b) and (c) indicate strong-coupling estimates. (d) Entanglement spectrum across the topological transition at $U/t=J_{\rm H}/t\approx 1.2$. (e,f) Spin–spin correlation function $\langle\mathbf{S}_{\gamma,i}\cdot\mathbf{S}_{\gamma^{\prime},j}\rangle$ as a function of distance $r$, shown on (e) log-log and (f) semi-log scales for $U/t=J_{\rm H}/t=0.8$ and 1.6, respectively.
@@
The relationship between spin gap and string order in the Hubbard ladder is more subtle. Figure 5(b) plots $\Delta_{\rm s}/t$ as a function of $J_{\rm H}/t$ for various $U/t$. A finite spin gap exists broadly, disappearing only at small $U/t$ and $J_{\rm H}/t$. A strong-coupling analysis yields an effective exchange $J_{\rm eff}=4t^{2}/(U+J_{\rm H}/2)$, predicting the asymptotic behavior $\Delta_{\rm s}/t=\{4t/(U+J_{\rm H}/2)\}\Delta_{\rm s}^{S=1}$. Remarkably, a finite gap persists even at $U/t=0$ for $J_{\rm H}/t\gtrsim 4$, reflecting an effective repulsion generated purely by strong Hund’s coupling [Fig. 5(c)]. Here, the gap asymptotically follows $\Delta_{\rm s}/t=(8t/J_{\rm H})\Delta_{\rm s}^{S=1}$.
Overlaying the finite spin-gap region onto Fig. 3, we find a perfect match with the finite string-order region. This nontrivial correspondence strongly supports the identification of this regime as a genuine Haldane phase.
@@
Finally, we examine real-space spin correlations to further distinguish the gapped Haldane phase from gapless trivial states. Figures LABEL:fig:spin_corr(e,f) present intra- and inter-leg spin correlations $\langle\mathbf{S}{\gamma,i}\cdot\mathbf{S}{\gamma^{\prime},j}\rangle$ as functions of distance $|i-j|$. In the Haldane phase ($U/t=J_{\rm H}/t=1.6$), the correlations decay exponentially as $\langle\mathbf{S}_{\gamma,i}\cdot\mathbf{S}_{\gamma^{\prime},j}\rangle\sim\exp(-|i-j|/\xi)$ ($\xi>0$), whereas in the gapless regime ($U/t=J_{\rm H}/t=0.8$), correlations show power-law decay, signaling criticality. Notably, within the Haldane phase, intra- and inter-leg correlations rapidly coincide, validating the emergence of effective spin-1 degrees of freedom and reinforcing the interpretation of the ladder as an effective spin-1 chain at low energies.

### v2

@@ Abstract
Introduction.— Understanding the robustness of topological phases in the presence of competing interactions and additional degrees of freedom remains a central challenge in condensed matter physics. Symmetry-protected topological (SPT) phases, exemplified by the Haldane phase [1, 2] in spin-1 chains, exhibit hallmark features including a finite spin gap [3, 4], nonlocal string order [5, 6, 7], and fractionalized edge states [8, 9] protected by discrete symmetries [10, 11, 12, 13]. While these properties have been thoroughly established for idealized spin-only models, their stability in realistic materials featuring significant charge fluctuations and orbital degrees of freedom remains poorly understood and is a highly nontrivial issue [14].
@@
Model definitions and effective mapping.— We consider two models of coupled one-dimensional (1D) chains with interchain FM Hund’s coupling: a spin-$\frac{1}{2}$ Heisenberg ladder and a half-filled Hubbard ladder. The Heisenberg ladder serves as a spin-only reference system, while the latter incorporates charge degrees of freedom and their fluctuations.
@@
$H_{\mathrm{Hub}}=-t\sum_{\gamma=1,2}\sum_{i,\sigma}\left(c^{\dagger}_{\gamma,i,\sigma}c_{\gamma,i+1,\sigma}+\text{H.c.}\right)$
@@
where $\tilde{\mathbf{S}}_{i}=\mathbf{S}_{1,i}+\mathbf{S}_{2,i}$ is a spin-1 operator on rung $i$, and $L$ is the number of rungs. This emergent spin-1 mechanism differs from materials with intrinsic local spin-1 ions (e.g., Ni2+), yet at low energies the two are adiabatically connected and realize the same Haldane SPT phase, as elaborated in the Supplemental Material [35].
@@
whose asymptotic value $|\mathcal{O}_{\mathrm{str}}(-\infty,\infty)|$ serves as the string order parameter. In practice, when the order is robust, $|\mathcal{O}_{\mathrm{str}}(i,j)|$ rapidly reaches a flat plateau as |ij||i-j| increases on large open ladders; in this case we read off the plateau value from a fixed, sufficiently long ladder and take it as the order parameter. When the plateau is not unambiguous (e.g., near phase boundaries), we determine the order parameter by finite-size scaling of a central estimator evaluated on $L\times 2$ ladders (see the Supplemental Material for details [35]). For the spin-1 Heisenberg chain, this parameter takes the value $|\mathcal{O}_{\mathrm{str}}^{S=1}(-\infty,\infty)|\approx 0.3743$ [4, 36].
@@
Figure 2: String order in the Heisenberg ladder. (a) String correlation function $\mathcal{O}_{\mathrm{str}}(|i-j|)$ plotted as a function of distance |ij||i-j| for several values of $J_{\rm H}/J$. The dotted line indicates the string order parameter of the $S=1$ Heisenberg chain. (b) String order parameter $|\mathcal{O}_{\rm str}(-\infty,\infty)|$ as a function of $J_{\rm H}/J$. Inset: $|\mathcal{O}_{\rm str}(-\infty,\infty)|$ plotted against $\sqrt{J_{\rm H}/J}$ in the small $J_{\rm H}/J$ regime; the dashed line shows a fit near $J_{\rm H}/J=0$.
Haldane Phase in the Heisenberg Ladder.— We first investigate how the Heisenberg ladder converges to the effective spin-1 Heisenberg chain [Eq. (3)] as $J_{\rm H}/J$ increases. Figure 2(a) shows the string correlation function $|\mathcal{O}_{\rm str}(i,j)|$ versus distance |ij||i-j| for several values of $J_{\rm H}/J$, computed on a $193\times 2$ ladder. For all $J_{\rm H}/J$ values considered, the correlation function clearly saturates at large distances, confirming robust long-range string order.
In Fig. 2(b), we plot the extracted string order parameter as a function of $J_{\rm H}/J$. As $J_{\rm H}/J$ grows, the string order parameter continuously approaches the known spin-1 Heisenberg chain value, demonstrating that the ladder progressively realizes the effective spin-1 limit. As shown in the inset, the string order parameter follows the scaling form $|\mathcal{O}_{\mathrm{str}}(i,j)|\propto\sqrt{J_{\rm H}/J}$ in the small-$J_{\rm H}/J$ regime, indicating the immediate onset of the Haldane phase for any finite $J_{\rm H}>0$. Consistent with this, the spin gap opens simultaneously with the emergence of string order, as discussed in detail below.
@@
Figure 4: Topological string order in the Hubbard ladder. String correlation function $O_{\text{str}}(i,j)$ versus |ij||i-j| for (a) various $J_{H}/t$ at fixed $U/t=4$ and (b) various $U/t$ at fixed $J_{H}/t=4$. Log-log plots of $O_{\text{str}}(i,j)$ versus |ij||i-j| for (c) various interaction strengths with $U/t=J_{H}/t$, and (d) various $J_{H}/t$ at fixed $U/t=0$. String order parameter as a function of (e) $U/t=J_{H}/t$ and (f) $J_{H}/t$ at fixed $U/t=0$. Dotted lines mark the spin-1 Heisenberg chain value.
Topological Order in the Hubbard Ladder.— We now investigate how charge fluctuations influence topological order in the Hubbard ladder by examining the string correlation function $\mathcal{O}_{\rm str}(i,j)$. Figure 4(a) presents $|\mathcal{O}_{\rm str}(i,j)|$ as a function of distance |ij||i-j| for several values of $J_{\rm H}/t$ at fixed $U/t=4$, while Fig. 4(b) shows similar results for various $U/t$ at fixed $J_{\rm H}/t=4$, computed on a $97\times 2$ ladder. In both cases, the string correlation saturates at finite values at large distances, confirming robust long-range order. With increasing $U/t$ or $J_{\rm H}/t$, charge fluctuations are progressively suppressed, driving the system closer to the effective spin-1 regime. Correspondingly, the saturated value, i.e., the string order parameter, approaches the spin-1 Heisenberg chain value. Thus, the Hubbard ladder can realize a stable Haldane phase despite moderate charge fluctuations.
@@
Long-range string order appears even at $U/t=0$, driven solely by Hund’s coupling. Figure 4(d) shows log-log plots of the string correlation for various $J_{\rm H}/t$ values at zero on-site repulsion. A clear transition from power-law decay ($J_{\rm H}/t=3$) to finite long-range order ($J_{\rm H}/t\geq 6$) is observed, reflecting strong suppression of charge fluctuations (double occupancy) by Hund’s coupling alone. This insulating phase stabilized purely by Hund’s coupling can thus be viewed as a novel “Hund insulator”. The string order parameter, shown in Fig. 4(f), grows around $J_{\rm H}/t\approx 4.0$, approaching the spin-1 Heisenberg value at large $J_{\rm H}/t$, again following an approximately linear onset.
The resulting phase diagram, summarized in Fig. 3, reveals a clear boundary between the trivial Mott/Hund insulating phase and the SPT Haldane phase, identified explicitly by the string order parameter. Notably, the Haldane phase remains stable under substantial charge fluctuations, persisting up to a local charge variance of approximately $\delta n^{2}=2\langle n_{\gamma,i,\uparrow}n_{\gamma,i,\downarrow}\rangle\lesssim 0.4$. Furthermore, the phase boundary approaches $J_{\rm H}=0$ in the $U=\infty$ limit, indicating a smooth crossover from the Hubbard to the Heisenberg ladder in the strongly correlated regime. Although our model setup is different, the qualitative agreement with the topological transition line reported in Ref. [26] supports the broader conclusion that the Haldane phase is generally stable against moderate charge fluctuations.
Figure 5: (a) Spin gap $\Delta_{\rm s}/J$ for the Heisenberg ladder as a function of $J_{\rm H}/J$. The dashed line marks the spin gap of the spin-1 Heisenberg chain [Eq. (3)]. Inset: Enlarged view of the small-$J_{\mathrm{H}}/J$ region. (b) $\Delta_{\rm s}/t$ for the Hubbard ladder as a function of $J_{\rm H}/t$ for several values of $U/t$. (c) $\Delta_{\rm s}/t$ as a function of $J_{\rm H}/t$ at $U=0$. Dashed lines in (b) and (c) indicate strong-coupling estimates. (d) Entanglement spectrum across the topological transition at $U/t=J_{\rm H}/t\approx 1.2$. Numbers beside the lines denote entanglement level degeneracies. (e,f) Spin–spin correlation function $\langle\mathbf{S}_{\gamma,i}\cdot\mathbf{S}_{\gamma^{\prime},j}\rangle$ as a function of distance $r$, shown on (e) log-log and (f) semi-log scales for $U/t=J_{\rm H}/t=0.8$ and 1.6, respectively.
@@
The relationship between spin gap and string order in the Hubbard ladder is more subtle. Figure 5(b) plots $\Delta_{\rm s}/t$ as a function of $J_{\rm H}/t$ for various $U/t$. A finite spin gap exists broadly, disappearing only at small $U/t$ and $J_{\rm H}/t$. A strong-coupling analysis yields an effective exchange $J_{\rm eff}=4t^{2}/(U+J_{\rm H}/2)$, predicting the asymptotic behavior $\Delta_{\rm s}/t=\{4t/(U+J_{\rm H}/2)\}\Delta_{\rm s}^{S=1}$. Also, a finite gap persists even at $U/t=0$ for $J_{\rm H}/t\gtrsim 4$, reflecting an effective repulsion generated purely by strong Hund’s coupling [Fig. 5(c)]. Here, the gap asymptotically follows $\Delta_{\rm s}/t=(8t/J_{\rm H})\Delta_{\rm s}^{S=1}$. Details of the strong-coupling analysis are given in the Supplemental Material [35].
Overlaying the finite spin-gap region onto Fig. 3, we find a perfect match with the finite string-order region. This nontrivial correspondence strongly supports the identification of this regime as a genuine Haldane phase. We also find that slightly away from half filling both the nonlocal string order and the spin gap remain finite in the infinitesimal doping limit (see the Supplemental Material [35]).
@@
Finally, we examine real-space spin correlations to further distinguish the gapped Haldane phase from gapless trivial states. Figures 5(e,f) present intra- and inter-leg spin correlations $\langle\mathbf{S}_{\gamma,i}\cdot\mathbf{S}_{\gamma^{\prime},j}\rangle$ as functions of distance |ij||i-j|. In the Haldane phase ($U/t=J_{\rm H}/t=1.6$), the correlations decay exponentially as $\langle\mathbf{S}_{\gamma,i}\cdot\mathbf{S}_{\gamma^{\prime},j}\rangle\sim\exp(-|i-j|/\xi)$ ($\xi>0$), whereas in the gapless regime ($U/t=J_{\rm H}/t=0.8$), correlations show power-law decay, signaling criticality. Notably, within the Haldane phase, intra- and inter-leg correlations rapidly coincide, validating the emergence of effective spin-1 degrees of freedom and reinforcing the interpretation of the ladder as an effective spin-1 chain at low energies.
@@
The data used to generate the figures in this manuscript are available in Ref. [41].